# MGDepth: Motion-Guided Cost Volume For Self-Supervised Monocular Depth In Dynamic Scenarios

### v1

@@ Abstract
Despite advancements in self-supervised monocular depth estimation, challenges persist in dynamic scenarios due to the dependence on assumptions about a static world. In this paper, we present MGDepth, a Motion-Guided Cost Volume Depth Net, to achieve precise depth estimation for both dynamic objects and static backgrounds, all while maintaining computational efficiency. To tackle the challenges posed by dynamic content, we incorporate optical flow and coarse monocular depth to create a novel static reference frame. This frame is then utilized to build a motion-guided cost volume in collaboration with the target frame. Additionally, to enhance the accuracy and resilience of the network structure, we introduce an attention-based depth net architecture to effectively integrate information from feature maps with varying resolutions. Compared to methods with similar computational costs, MGDepth achieves a significant reduction of approximately seven percent in root-mean-square error for self-supervised monocular depth estimation on the KITTI-2015 dataset.
keywords:
Unsupervised Learning, Depth Estimation, Ego-motion Estimation, Dynamic Scenario, Attention Mechanism.
[…]
[label1]organization=University Of Oxford, city=Oxford, postcode=OX1 2JD, country=United Kingdom
1 Introduction
@@
The role of vision-based depth estimation (VDE) has become increasingly important in computer vision due to its ability to understand the 3D geometry of a scene based on the 2D observation, which serves as the foundation for various high-level 3D tasks such as scene reconstruction niemeyer2022regnerf ; zhou2023dynpoint ; li2021hierarchical , object detection wang2021salient and navigation de2021deep . Moreover, VDE has enabled state-of-the-art applications ranging from autonomous driving you2019pseudo ; wang2019pseudo ; miao2023ds to augmented reality newcombe2011dtam ; mieloch2022overview ; noraky2019low .
Figure 1: Qualitative comparison on KITTI-2015. (a) shows the RGB image of the current frame. (b) displays the estimated depth generated by Monodepth2 trained with monocular mode. (c) presents the estimated depth of Monodepth2 trained with stereo mode. (d) demonstrates the estimated depth of MGDepth in monocular mode. Differences are masked out by circles.
@@
Recently, self-supervised depth estimation has emerged as a viable approach for training depth estimation methods, aiming to alleviate the dependency on extensive training data and reduce high computational demands. These methods learn depth maps from either monocular images bian2021unsupervised or stereo image pairs yao2018mvsnet . Despite the significant advancements made in self-supervised monocular vision-based depth estimation, a notable performance gap persists when comparing self-supervised monocular VDEs to self-supervised stereo VDEs. The disparity in performance can be mainly ascribed to the capability of stereo methods to utilize multiple views for constructing a feature volume, thereby incorporating a greater amount of 3D camera frustum information. While multi-frame monocular VDEs, as presented in watson2021temporal , have the ability to construct a feature volume based on adjacent frames, the presence of dynamic elements in these adjacent frames can potentially disrupt the construction of the feature volume.
@@
Taking the aforementioned concern into account, this paper strives to enhance the performance of multi-frame monocular VDEs by integrating motion information during the inference process and adeptly managing the influence of dynamic objects in the construction of the feature volume. Experimental results indicate that the proposed multi-frame monocular VDEs can surpass even previous monocular VDEs that directly incorporate stereo information during the training process, as demonstrated in the examples provided in Fig. 1.
@@
Specifically, in light of recent advancements in optical flow estimation and its successful application in motion detection and estimation tasks, we propose MGDepth. MGDepth is a self-supervised monocular vision-based depth estimation (VDE) system that leverages motion information through a motion-guided cost volume constructed with an attention mechanism. The attention mechanism is chosen due to its demonstrated outstanding performance in representation learning and effective fusion of diverse information. Our main contributions are summarized as follows:
  • 1.
@@
The structure of this paper unfolds as subsequent sections detail various aspects. Section II offers a concise overview of related studies concerning monocular depth estimation techniques. Section III delves into the intricacies of the proposed MGDepth. Findings from experiments and a thorough model analysis are outlined in Section IV. Section V wraps up with the conclusions drawn from the presented research.
2 Related Works
2.1 Monocular Depth Estimation
Self-supervised techniques for monocular depth prediction have gained notable traction due to their independence from labeled data and their versatility across various scenarios. Broadly, monocular VDEs fall into two categories: one relies solely on the present frame for depth estimation, while the other employs multiple neighboring frames to achieve depth prediction.
2.1.1 Single-frame Monocular Depth Estimation
In the initial category, Monodepth zhou2017unsupervised emerges as a leading benchmark, showcasing reliable results. The model features two networks: one focused on pose estimation and the other on depth estimation. During the optimization process, these networks synergize by utilizing the warping connection between depth and image transformations to construct a photometric loss function. Building upon this framework, Monodepth2 godard2019digging introduces a minimal projection error approach to tackle occlusion and minimize visual artifacts through full-resolution multiscale sampling. In addition to employing traditional photometric loss, FeatDepth shu2020feature introduces a feature-space-based metric loss to more effectively regulate the warping relationship between target and source frames. Moreover, DevNet zhou2022devnet enhances its performance by integrating 3D geometric consistency among neighboring frames.
2.1.2 Multi-frame Monocular Depth Estimation
@@
The premise of the second category rests on the notion that integrating temporal information during inference—by employing multiple neighboring frames as inputs—can enhance the accuracy of the final depth estimation. Initially, this is accomplished by employing test-time refinement techniques casser2019depth ; chen2019self ; mccraith2020monocular ; kuznietsov2021comoda , along with recurrent neural networks (RNNs) as evidenced in studies like patil2020don ; wang2019recurrent ; zhang2019exploiting . The test-time refinement method adopts a monocular strategy to leverage temporal data during testing, whereas the recurrent neural network integrates with a monocular depth estimation network to analyze continuous frame sequences. Nonetheless, models utilizing recurrent neural networks often entail high computational costs and lack a distinct geometric inference approach. Recently, Manydepth watson2021temporal and MonoRec wimbauer2021monorec have made notable advancements in performance and real-time efficiency by incorporating cost volumes from stereo-matching tasks for geometric-based reasoning gu2020cascade . These models rely on a photometric loss function, where temporally neighboring frames are mapped onto the current image plane using predetermined depth bins. Within the cost volume framework, the inferred depth with the minimal value corresponds closely to the actual depth.
Nevertheless, these approaches are grounded in static assumptions regarding scenarios and struggle with dynamic foreground elements. To tackle this limitation, we introduce MGDepth, a technique adept at managing dynamic foreground by integrating temporal data into the cost volume and implementing a motion-guided photometric loss function.
Figure 2: Illustration for the structure of MGDepth. During Stage 1, target and reference frame $I_{t}$ and $I_{r}$ are processed Flow Net, Pose Net, and Depth Net to generate optical flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$. During Stage 2, the outputs are used to generate the motion-guided cost volume $\mathcal{V}_{m}$. Finally, the motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used by Attention-Based Depth Net to produce the refined depth $D_{t}$.
2.2 Self-supervised VDE for Dynamic Objects
Due to the distinctive characteristics of dynamic objects, researchers have pursued differential treatment of these objects during both the training and inference processes in self-supervised vision-based depth estimation (VDE) methods. In klingner2020self , dynamic objects are deliberately excluded from consideration to prevent their influence on the optimization process and maintain the accuracy of the depth estimation. In feng2022disentangling ; ranftl2016dense , dynamic objects are initially segmented and subsequently treated differently in the photometric loss function. This approach allows for the incorporation of dynamic objects into the depth learning process, enabling their meaningful contribution rather than direct exclusion. Moreover, in lee2021learning ; lee2021attentive ; li2021unsupervised , an effort is made to predict object-level motion information and leverage it to provide enhanced constraints for self-supervised depth learning. However, these methods either suffer from high computational demands or exhibit limited performance improvement. Our proposed method, MGDepth, addresses these limitations by constructing a motion-guided cost volume that integrates predicted optical flow information through an attention mechanism.
3 Method
3.1 Overview
The primary structure of MGDepth is based on previous self-supervised monocular depth estimation algorithms, such as Monodepth godard2017unsupervised and Manydepth watson2021temporal , with the goal of estimating depth from monocular video recorded by an unknown camera. However, existing methods face difficulties in accurately estimating depth for moving objects lee2021learning ; li2021unsupervised . We propose the following innovative strategies to achieve accurate monocular depth estimation in complex dynamic scenarios. (1) We make use of the optical flow information and prior depth information to construct a new static reference frame. (2) Based on the new static reference frame, the original reference frame, and the new target frame, we construct a motion-guided cost volume. (3) We introduce an attention-based depth net structure, specifically designed to proficiently amalgamate feature maps with varying resolutions derived from a high-resolution network.
3.2 Framework
@@
The structure of MGDpth is demonstrated in Fig. 2. At the outset, MGDpth employs a pretrained optical flow network denoted by $\theta_{f}$ and a pose network denoted by $\theta_{p}$ to predict the optical flow $f_{r\leftarrow t}$ and the transformation matrix $[R|\bm{t}]_{rt}$ from the target frame to reference frame. Concurrently, a prior depth network denoted by $\theta_{cd}$ is utilized to estimate the coarse depth map $D_{t}^{c}$ for the target frame. Subsequently, the target frame denoted as $I_{t}$, the reference frame represented by $I_{r}$, the coarse depth map denoted as $D_{t}$, the transformation matrix $[R|\bm{t}]_{rt}$, and the optical flow $f_{r\leftarrow t}$ are utilized collaboratively to construct a motion cost volume denoted as $\mathcal{V}_{m}$. The motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used in unison as inputs to the attention-based depth network $\theta_{ad}$ with the objective of predicting the final depth map $D_{t}$. Subsequently, this predicted depth map $D_{t}$ is utilized in the construction of the ultimate motion-guided photometric loss denoted as $\mathcal{L}_{mp}$.
@@
3.3 New Static Reference Frame
@@
where $p_{t}$ are pixels in the frame $I_{t}$; $K_{r}\&K_{t}$ are intrinsic matrix for frames $I_{r}\&I_{t}$. As illustrated in Fig. 3, in static scenarios, the static flow $f^{\prime}_{r\leftarrow t}$ aligns with the real optical flow $f_{r\leftarrow t}$. In scenarios involving moving objects, the real optical flow $f_{r\leftarrow t}$ can be decomposed into static optical flow $f^{\prime}_{r\leftarrow t}$ and dynamic optical flow $f^{d}_{r\leftarrow t}$. Then the dynamic flow can be calculated as $f^{d}_{r\leftarrow t}=f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}$. The motion mask could be generated as follow:
@@
where $\epsilon$ is the threshold for distinguishing moving parts. The motion pattern observed in the reference frame can be utilized to create a corresponding mask in the target frame. This is achieved through the generation of a static motion mask $\mathcal{M}_{ms}^{r}$, utilizing the static optical flow $f^{\prime}_{r\leftarrow t}$. Simultaneously, the dynamic motion mask $\mathcal{M}_{md}^{r}$ is derived with the assistance of $f_{r\leftarrow t}$. The composite motion mask for the reference frame is obtained by combining these two masks, denoted as $\mathcal{M}_{m}^{r}=\mathcal{M}_{ms}^{r}\cup\mathcal{M}_{md}^{r}$. The rationale for avoiding the direct computation of mask $\mathcal{M}_{m}^{r}$ using $f^{\prime}_{t\leftarrow r}$ lies in the necessity of inferring the depth of the target frame $D_{r}$. This procedure would subsequently lead to a rise in computational demands for both the training and inference processes.
@@
With $\mathcal{M}_{m}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
$\hat{I_{r}}=\begin{cases}I_{r}&p_{i}\notin\mathcal{M}_{m};\\ \phi^{\prime}(I_{t},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt})&p_{i}\in\mathcal{M}_{m}.\\ \end{cases}$ (4)
@@
It is important to note that $\phi^{\prime}(.)$ differs from $\phi(.)$, as $\phi(.)$ is utilized to generate current frame information via sampling based on pixel correspondence between the current frame and the adjacent frame, whereas $\phi^{\prime}(.)$ generates adjacent frame by shifting forward pixels of the current frame. Please note that $\hat{I_{r}}$ is detached to eliminate its impact on the gradient calculation of backpropagation process.
3.4 Motion-Guided Cost Volume
Figure 4: Illustration for the construction of Final Volume. Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$. This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
@@ 3.4 Motion-Guided Cost Volume
In constructing the motion-guided cost volume $\mathcal{V}_{m}$ for target frame $I_{t}$, a set of parallel planes that are perpendicular to the optic axis of $I_{t}$ is defined, based on the depth assumptions $\mathcal{D}=\{d_{k};k=1,...,M\}$, where $M$ represents the number of planes. The feature extractor $\theta_{fe}$ is utilized to generate the feature maps $\mathcal{F}_{t}$, $\mathcal{F}_{r}$ and $\hat{\mathcal{F}_{r}}$ of $I_{t}$, $I_{r}$ and $\hat{I_{r}}$, respectively. With the aid of $\mathcal{D}$ and $[R|\bm{t}]_{rt}$, a set of wrapped feature maps $\{\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}},\mathcal{F}_{t\leftarrow r}^{d_{k}};d_{k}\in\mathcal{D}\}$ are generated by warping $\hat{\mathcal{F}_{r}}$ and $\mathcal{F}_{r}$. Considering $\mathcal{V}_{m}=\{\mathcal{P}_{k},k=1,...,M\}$, it could be written as:
$\mathcal{P}_{k}=\alpha\frac{\sum_{i=1}^{N}|\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}}-\mathcal{F}_{t}|}{N}+(1-\alpha)\frac{\sum_{i=1}^{N}|\mathcal{F}_{t\leftarrow r}^{d_{k}}-\mathcal{F}_{t}|}{N}.$ (5)
Figure 5: Illustration for the construction of Final Volume. Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$. This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
Here, $N$ denotes the number of reference frames used in construction, while in our experiment, we typically use a single adjacent frame as a reference frame. $\alpha$ denotes the parameters used to balance the influence of new static reference frame $\hat{I_{r}}$. The main reason for keeping the first item in the motion-guided cost volume is that at the early stage of the training process, the prior depth is not accurate enough which makes the $f^{\prime}_{r\leftarrow t}$ noisy. With the progress of training, $\alpha$ will gradually decrease to improve the influence of the new static reference frame in the construction of $\mathcal{V}_{m}$.
@@
Instead of directly taking the volume $\mathcal{V}_{m}$ into the later depth network, we propose to construct the final feature volume $\mathcal{V}_{f}$ by combining the $\mathcal{V}_{m}$ with the $\mathcal{F}_{t}$ though channel attention. The process of constructing the final feature volume $\mathcal{V}_{f}$ is presented in Fig. 5.
Figure 6: Illustration for the structure of Attention-based Depth Net. To effectively leverage feature maps of varying resolutions, we employ efficient channel attention to concatenate high-resolution feature maps with low-resolution feature maps and the low-resolution disparity map. This resultant feature map is subsequently utilized for disparity prediction within the current branch and is propagated to the subsequent branch.
3.5 Attention-Based Depth Network
The High-Resolution Network (HRNet) wang2020deep is well-regarded for its ability to preserve a high level of detail in input images. Although it is widely used in various computer vision tasks, its performance in Vision Depth Estimation (VDE) has not been extensively explored. To enhance the accuracy and efficiency of VDE, we integrate an efficient channel attention module into the design of the depth decoder network of the HRNet-based depth network.
The HRNet is composed of multiple branches, denoted as $B$, each generating $S=b,...,4$ features with resolutions of $(\frac{H}{2^{b-1}},\frac{W}{2^{b-1}})$. Following the framework of Monodepth2, we utilize all branches to produce multi-scale disparity maps denoted as $D^{b}$, and apply the same loss function to each branch.

### v2

# Manydepth2: Motion-Aware Self-Supervised Monocular Depth Estimation in Dynamic Scenes

@@ Abstract
Despite advancements in self-supervised monocular depth estimation, challenges persist in dynamic scenarios due to the dependence on assumptions about a static world. In this paper, we present Manydepth2, a Motion-Guided Cost Volume Depth Net, to achieve precise depth estimation for both dynamic objects and static backgrounds, all while maintaining computational efficiency. To tackle the challenges posed by dynamic content, we incorporate optical flow and coarse monocular depth to create a novel static reference frame. This frame is then utilized to build a motion-guided cost volume in collaboration with the target frame. Additionally, to enhance the accuracy and resilience of the network structure, we introduce an attention-based depth net architecture to effectively integrate information from feature maps with varying resolutions. Compared to methods with similar computational costs, Manydepth2 achieves a significant reduction of approximately five percent in root-mean-square error for self-supervised monocular depth estimation on the KITTI-2015 dataset. The code could found here .
Index Terms:
Monocular Depth Estimation, Self-Supervised, Optical Flow.
11footnotetext: KZ, JB, JZ, QX, NT and AM are with the University of Oxford. Contact Email: zhouk777@mit.edu
I Introduction
@@
The role of vision-based depth estimation (VDE) has become increasingly important in computer vision due to its ability to understand the 3D geometry of a scene based on 2D observation, which serves as the foundation for various high-level 3D tasks, such as scene reconstruction [1], object detection [2] and navigation [3]. Moreover, VDE has enabled state-of-the-art applications ranging from autonomous driving [4] to augmented reality [4].
@@
Recently, self-supervised depth estimation has emerged as a viable approach for training depth estimation methods, aiming to alleviate the dependency on extensive training data and reduce high computational demands. These methods learn depth maps from either monocular images [5] or stereo image pairs [6]. Despite the significant advancements made in self-supervised monocular vision-based depth estimation, a notable performance gap persists when comparing self-supervised monocular VDEs to self-supervised stereo VDEs. The disparity in performance can be mainly ascribed to the capability of stereo methods to utilize multiple views for constructing a feature volume, thereby incorporating a greater amount of 3D camera frustum information. While multi-frame monocular VDEs, as presented in [7], have the ability to construct a feature volume based on adjacent frames, the presence of dynamic elements in these adjacent frames can potentially disrupt the construction of the feature volume.
Figure 1: Qualitative comparison on Cityscapes. The first row presents the RGB images. The second row shows the error maps of the depth estimates produced by ManyDepth. The third row displays the error maps of the depth estimates generated by Manydepth2.
@@
Taking the aforementioned concern into account, this paper strives to enhance the performance of multi-frame monocular VDEs by integrating motion information during the inference process and adeptly managing the influence of dynamic objects in the construction of the feature volume. Experimental results indicate that our method could effectively handle dynamic objects, as demonstrated in Fig. 1.
@@
Specifically, in light of recent advancements in optical flow estimation and its successful application in motion detection and estimation tasks, we propose Manydepth2. Manydepth2 is a self-supervised monocular vision-based depth estimation system that leverages motion information through a motion-guided cost volume constructed with an attention mechanism. The attention mechanism is chosen due to its demonstrated outstanding performance in representation learning and effective fusion of diverse information. Our main contributions are summarized as follows:
Figure 2: Illustration for the structure of Manydepth2. During Stage 1, target and reference frame $I_{t}$ and $I_{r}$ are processed Flow Net, Pose Net, and Depth Net to generate optical flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$. During Stage 2, the outputs are used to generate the motion-guided cost volume $\mathcal{V}_{m}$. Finally, the motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used by Attention-Based Depth Net to produce the refined depth $D_{t}$.
@@
II Related Works
II-A Monocular Depth Estimation
@@
Self-supervised techniques for monocular depth prediction have gained notable traction due to their independence from labeled data and their versatility across various scenarios. Broadly, monocular VDEs fall into two categories: one relies on the present frame for depth estimation, while the other employs multiple frames to achieve depth prediction. Single-frame Monocular Depth Estimation: Within the realm of depth estimation methodologies, the initial category has witnessed the emergence of notable frameworks that have set benchmarks in terms of reliability and performance. A prime example from this category is Monodepth, as highlighted in the study by [8]. This model distinguishes itself with a bifurcated network architecture: one part is dedicated explicitly to pose estimation, while another part focuses on depth estimation tasks. A pivotal aspect of the Monodepth architecture lies in the collaborative synergy between these two networks. They operate by leveraging the warping connection intrinsic to depth and image transformations. This collaboration results in creating a photometric loss function, which is a key part of why it works so well. Further elevating the landscape of depth estimation methodologies, DevNet, in [9], takes a comprehensive approach by embedding 3D geometric consistency. Multi-frame Monocular Depth Estimation: The premise of the second category rests on the notion that integrating temporal information during inference—by employing multiple neighboring frames as inputs—can enhance the accuracy of the final depth estimation. Initially, this is accomplished by employing test-time refinement techniques [10], along with recurrent neural networks (RNNs) as evidenced in studies like [11]. The test-time refinement method adopts a monocular strategy to leverage temporal data during testing, whereas the recurrent neural network integrates with a monocular depth estimation network to analyze continuous frame sequences. Nonetheless, models utilizing recurrent neural networks often entail high computational costs and lack a distinct geometric inference approach. Recently, Manydepth [7] and MonoRec [12] have made notable advancements in performance and real-time efficiency by incorporating cost volumes from stereo-matching tasks for geometric-based reasoning [13]. These models rely on a photometric loss function, where temporally neighboring frames are mapped onto the current image plane using predetermined depth bins. Within the cost volume framework, the inferred depth with the minimal value corresponds closely to the actual depth. Nevertheless, these approaches are grounded in static assumptions regarding scenarios and struggle with dynamic foreground elements. To tackle this limitation, we introduce Manydepth2, a technique adept at managing dynamic foreground by integrating temporal data into the cost volume and implementing a motion-guided photometric loss function.
II-B Self-supervised VDE for Dynamic Objects
Owing to the unique attributes of dynamic objects, there has been a focused effort among researchers to handle these objects distinctively throughout both the training and inference stages within self-supervised vision-based depth estimation (VDE) techniques. In [14], a deliberate decision was made to exclude dynamic objects from the analysis. This exclusion was aimed at preventing any potential interference with the optimization process, thereby ensuring the precision and accuracy of the depth estimation results. On the other hand, in [15], a more nuanced approach was adopted. Dynamic objects were first identified and segmented from the scene. Subsequent to this segmentation, these objects were treated distinctively within the framework of the photometric loss function. Such a strategy facilitates the integration of dynamic objects into the depth learning mechanism, ensuring that they contribute meaningfully without merely being sidelined or excluded. Furthermore, recent research efforts, as seen in [16], have ventured into predicting motion information at the object level. This predictive capability is then harnessed to establish more refined constraints, thereby enhancing the efficacy of self-supervised depth learning methodologies. Despite these advancements, it’s noteworthy that many of these approaches either grapple with intensive computational requirements or demonstrate only marginal enhancements in performance metrics. Recognizing these challenges, our novel approach, termed Manydepth2, seeks to overcome these constraints. We introduce a motion-guided cost volume framework that seamlessly integrates forecasted optical flow insights, all achieved through a sophisticated attention mechanism.
III Methodology
III-A Overview
Firstly, as a foundational step, we utilize the potential of both optical flow information and pre-training depth data. By combining these insights, we construct a novel static reference frame. This innovative approach ensures that the depth estimation process remains robust and relevant, even when confronted with dynamic elements within the scene. With this static reference frame, our methodology further integrates the novel static reference frame alongside the original target frame. Combining these frames leads to creating a carefully designed motion-guided cost volume. This volume is crucial, improving the depth estimation by considering the detailed movement and space relationships in the scene. Finally, We are introducing a depth net structure that utilizes attention mechanisms. This structure is meticulously crafted to effectively combine feature maps that come from different resolutions, all originating from a high-resolution network.
III-B Framework
@@
The structure of Manydepth2 is demonstrated in Fig. 2. At the outset, Manydepth2 employs a trained optical flow network denoted by $\theta_{f}$ and a pose network denoted by $\theta_{p}$ to predict the optical flow $f_{r\leftarrow t}$ and the transformation matrix $[R|\bm{t}]_{rt}$ from the target frame to reference frame. Concurrently, a prior depth network denoted by $\theta_{cd}$ is utilized to estimate the coarse depth map $D_{t}^{c}$ for the target frame. Subsequently, the target frame denoted as $I_{t}$, the reference frame represented by $I_{r}$, the coarse depth map denoted as $D_{t}$, the transformation matrix $[R|\bm{t}]_{rt}$, and the optical flow $f_{r\leftarrow t}$ are utilized collaboratively to construct a motion cost volume denoted as $\mathcal{V}_{m}$. The motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used in unison as inputs to the attention-based depth network $\theta_{ad}$ with the objective of predicting the final depth map $D_{t}$. Subsequently, this predicted depth map $D_{t}$ is utilized in the construction of the ultimate motion-guided photometric loss denoted as $\mathcal{L}_{mp}$.
@@
III-C New Static Reference Frame (Rewrite and Check)
@@
where $p_{t}$ are pixels in the frame $I_{t}$; $K_{r}\&K_{t}$ are intrinsic matrix for frames $I_{r}\&I_{t}$. As illustrated in Fig. 3, in static scenarios, the static flow $f^{\prime}_{r\leftarrow t}$ aligns with the real optical flow $f_{r\leftarrow t}$. In scenarios involving moving objects, the real optical flow $f_{r\leftarrow t}$ can be decomposed into static optical flow $f^{\prime}_{r\leftarrow t}$ and dynamic optical flow $f^{d}_{r\leftarrow t}$. Then the dynamic flow can be calculated as $f^{d}_{r\leftarrow t}=f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}$.
Figure 4: Qualitative results on KITTI Dataset. The initial row displays the RGB images of target frames where depth information has been estimated. The subsequent rows, specifically the second, third, and fourth, showcase depth maps generated by Monodepth2 trained with stereo techniques, ManyDepth, and Manydepth2, respectively.
The motion mask could be generated as following:
@@
where $\epsilon$ is the threshold for distinguishing moving parts. The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}m^{r}$ in the reference mask, utilizing $f{r\leftarrow t}$. With $\mathcal{M}_{m}^{r}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
$\hat{I_{r}}=\begin{cases}I_{r}&p_{i}\notin\mathcal{M}_{m}^{r};\\ \phi^{\prime}(I_{t},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt})&p_{i}\in\mathcal{M}_{m}^{r}.\\ \end{cases}$ (4)
@@
It is important to note that $\phi^{\prime}(.)$ differs from $\phi(.)$, as $\phi(.)$ is utilized to generate current frame information via sampling based on pixel correspondence between the current frame and the adjacent frame, whereas $\phi^{\prime}(.)$ generates adjacent frame by shifting forward pixels of the current frame.
III-D Motion-Guided Cost Volume
@@
In constructing the motion-guided cost volume $\mathcal{V}_{m}$ for target frame $I_{t}$, a set of parallel planes that are perpendicular to the optic axis of $I_{t}$ is defined, based on the depth assumptions $\mathcal{D}=\{d_{k};k=1,...,M\}$, where $M$ represents the number of planes. The feature extractor $\theta_{fe}$ is utilized to generate the feature maps $\mathcal{F}_{t}$ and $\hat{\mathcal{F}_{r}}$ of $I_{t}$ and $\alpha\hat{I_{r}}+(1-\alpha)I_{r}$, respectively. With the aid of $\mathcal{D}$ and $[R|\bm{t}]_{rt}$, a set of wrapped feature maps $\{\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}};d_{k}\in\mathcal{D}\}$ are generated by warping $\hat{\mathcal{F}_{r}}$. Considering $\mathcal{V}_{m}=\{\mathcal{P}_{k},k=1,...,M\}$, it could be written as:
$\mathcal{P}_{k}=\frac{\sum_{i=1}^{N}|\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}}-\mathcal{F}_{t}|}{N}$ (5)
@@
Here, $N$ denotes the number of reference frames used in construction, while in our experiment. $\alpha$ denotes the parameters used to balance the influence of new static reference frame $\hat{I_{r}}$. Instead of directly taking the volume $\mathcal{V}_{m}$ into the later depth network, we propose to construct the final feature volume $\mathcal{V}_{f}$ by combining the $\mathcal{V}_{m}$ with the $\mathcal{F}_{t}$ though channel attention. The process of constructing the final feature volume $\mathcal{V}_{f}$ is presented in Fig. 2.
III-E Attention-Based Depth Network